A block slides down an inclined surface ofinclination 30 degree with the horizontal Starting from rest it covers 8m in the first two seconds .Find the coefficient of kinetic friction between the two

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

Force acting along the inclination is = mgsinΘ  - μmgcosΘ

so acceleration a =(mgsinΘ  - μmgcosΘ)/m

                         =gsinΘ  - μgcosΘ

S =ut +1/2 a t2  

8=0+1/2 (gsinΘ  - μgcosΘ)*4

or 4=10*1/2  - μ*10*√3/2

  μ=1/5√3


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin